HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

TOWARDS A HYDROGEN MOLECULAR ION CLOCK

\underline{M. BEYER}, Department of Physics and Astronomy, VU Amsterdam, Amsterdam, The Netherlands.

The molecular hydrogen ion is the simplest of all molecules and as such is an important system for the development of molecular quantum mechanics. The rovibrational energy-level structure of this one-electron system can be calculated extremely precisely by quantum-chemical methods which include the determination of relativistic and QED effects. By comparison with the results of precision measurements of rovibronic intervals, fundamental constants, such as the proton-to-electron mass ratio, can be determined.

I will give an overview of our work towards building a novel molecular clock based on weakly bound states of the molecular hydrogen ion. Various laser sources in the visible and (vacuum) ultraviolet wavelength range are used to selectively produce H_{2}^{+} in the desired weakly bound states, and to precisely measure its ’clock’ frequency in the microwave range.

Together, the theoretical framework and results of the clock measurements form a unique system that will shed light on several intriguing pending issues in physics, ranging from the (still not understood) role of nuclear-spin-induced symmetry breaking within the framework of quantum electrodynamics, to the possible existence of time-varying fundamental constants and hypothesized spin-dependent fifth forces.

p-number: p253  Submitted on Tue, 27 Jun 2023 21:34:32 +0200